# Large Language Models for cross-language code clone detection

### v1

@@ 3.5. Methodology
We overview the experimental methodology that is employed for assessing LLMs as well as classification models for the task of cross-lingual code clone detection. Figure 1 illustrates the process, highlighting the input datasets, the models as well as the expected outputs.
@@
Step 2 : Analyze the following two code snippets and determine if they are code clones. The function of the first code is step 1 output and the second is {step 1 output}. Please answer ‘yes’ if the code snippets are clones, regardless of the programming language, or ‘no’ if they are not. yes/no
@@
Cross-lingual code clone detection as a Classification task. We investigate, under the same conditions as LLMs, the potential for traditional classification mechanisms based on machine learning models. To that end, we follow the steps used by Keller et al. (Keller et al., 2022) in their learning-based approach for code clone detection. In their approach a classifier is trained based on sample sets of clone and non-clone pairs where each pair is represented by the absolute difference between the embedding vectors of each code fragment in the pair. Compared to other concatenation techniques that were assessed, their work conclude that differencing achieved the best results. In our case, however, the representations are yielded Embedding models that are provided alongside the LLM infrastructure of OpenAI. We use Text-embedding-Ada-002  [^5] https://openai.com/index/new-and-improved-embedding-model/ embedding model, which was trained on massive code and text data to yield representations that maximally capture semantics and can serve for similarity analysis. Finally, we train a binary classifier based on the yielded representations and using two basic learners: k Nearest Neighbors (k-NN) and Support Vector Machines (SVM).
Beyond the learning-based classification approach, we also investigate a straightforward approach where we compute the similarity between two fragments in a cross-lingual pair represented in a single representation space. To that end we apply cosine similarity metric on the Ada embeddings and check for the optimal threshold value to achieve the highest performance in terms of clone pair identification.
@@ 4. Evaluation
Simple prompt Non-clone 0.92 0.94 0.93 0.97 0.59 0.73
Clone 0.94 0.92 0.93 0.31 0.91 0.47
Similar line Non-clone 0.92 0.96 0.94 0.92 0.61 0.73
Clone 0.96 0.92 0.94 0.41 0.83 0.55
Reasoning Non-clone 0.96 0.95 0.96 0.94 0.62 0.75
Clone 0.95 0.96 0.96 0.42 0.88 0.57
Integrate Non-clone 0.96 0.90 0.93 0.96 0.58 0.72
Clone 0.89 0.96 0.92 0.30 0.88 0.45
Separate code Non-clone 0.99 0.89 0.94 0.99 0.54 0.70
Clone 0.87 0.99 0.93 0.15 0.97 0.26
Separate explanations Non-clone 0.99 0.97 0.98 0.99 0.58 0.73
Clone 0.97 0.99 0.98 0.30 0.97 0.45
@@
XLCoST CodeNet LLM Clone type Recall Precision F1-score Recall Precision F1-score Falcon-Instruct-7B Non-clone 0.08 0.37 0.13 0.07 0.51 0.12 Clone 0.87 0.49 0.62 0.94 0.5 0.65 StarChat-Beta-16B Non-clone 0.95 0.49 0.65 0.98 0.51 0.67 Clone 0.03 0.38 0.05 0.06 0.71 0.10 LLaMA2-Chat-7B Non-clone 0.00 1.00 0.00 0.04 0.77 0.07 Clone 1.00 0.50 0.67 0.99 0.51 0.67 GPT-3.5-Turbo Non-clone 0.92 0.94 0.93 0.97 0.59 0.73 Clone 0.94 0.92 0.93 0.31 0.91 0.47
@@
Goal. We aim to fill a gap in the literature with a comprehensive analysis of the effectiveness of LLMs on the task of cross-lingual code clone detection. In particular, we investigate the impact of state-of-art prompt engineering strategy on the performance of the LLMs: we thus experiment with Zero-shot and Chain-of-Thought (CoT)-based prompts. Depending on the desired output formats, the designed prompts fall into two categories:
@@
Experiments. We evaluated the four LLMs considered in this study (cf. Section 3.2), namely GPT-3.5-Turbo, Starchat-$\beta$-16B, Falcon-7B-Instruct, and LLAMA2-Chat-7B, on the randomly constructed subsets Section 3.3) from XLCoST and CodeNet. We use the ChatGPT API to prompt the GPT-3.5-Turbo LLM and parse the outputs to compute the precision, recall, and F1 scores finally. For prompts leading to outputs with a similarity score, we analyzed the LLMs’ outputs across a range of threshold values for code clone decisions.
@@
Results. Our experiments confirmed that many LLMs actually fail to support the Chain-of-Thought prompting strategy, confirming a claim by Dou et al. (Dou et al., 2023). Consequently, to analyze the performance of LLMs across different prompts, we focus on GPT-3.5-Turbo, which produced results when instructed with all our designed prompts. Table 2 summarizes the results for this LLM across all the ”yes/no” output prompts. On the XLCoST dataset, the CoT-based ”Separate explanations” prompt enables the model to achieve the highest F1 score (98% for non-clones and clones). However, on the CodeNet dataset, the same prompt achieves the second-lowest score (%), just ahead of the ”Separate code” prompts. This discrepancy in performance is due to the difference in complexity between the two datasets. Indeed, XLCoST is mainly composed of source code implementing simple problems such as “Compute modulus division by a power” while CodeNet is built based on coding challenges of higher complexity such as “Given a string, find the minimum number of swaps needed to rearrange the characters into a palindrome”. Furthermore, some LLMs exhibit a clear difference in their ability to detect code clones compared to non-clones. For instance, on the CodeNet dataset, the GPT-3.5-turbo model shows a 26 percentage point variation in the f1 score (73% on non-clone vs. 47% on clones). On the same dataset, we also observe a high variation of 53 for the ‘Falcon-Instruct-7B’ (12% on non-clones vs. 65% on clones) model. Manual analysis revealed a tendency for this model to classify nearly all input code snippet pairs as clones, resulting in an inflated number of false positives.
@@
[RQ-1.1] ☛ In general, for simple and recurrent programming tasks, LLMs show a high ability to detect cross-lingual clones (e.g., XLCoST samples). However, when applied to code samples of challenging programming tasks (e.g., CodeNet), LLMs performance drop by ${}30\%$ on average in terms of F1 score. This result is insightful as it calls for a different perspective in benchmarking cross-lingual code clone detection approaches. Indeed, while XLCost has been built for such purpose, CodeNet appears to offer a more realistic setting, in terms of complexity for assessment.
@@
The ”Simple prompt” exhibits a relatively consistent performance across both datasets, suggesting its potential for broader applicability. A detailed performance comparison of all four LLMs using this prompt is presented in Table 3. Notably, the GPT-3.5-turbo model outperformed other LLMs in this analysis.
@@
[RQ-1.2] ☛ Compared to Zero-Short prompting, which is mainly used in the literature, Chain of Thought prompting improves the performance of LLMs by 5-6 percentage points in terms of F1. It is further noteworthy that GPT-3.5-Turbo substantially outperforms the other studied LLMs in the task of cross-lingual code clone detection.
@@
In Figure 2, we present the result of the GPT-3.5-turbo LLM with the ”Code Similarity” prompt, designed to generate a similarity score between code snippets (0-10 scale) The results show that with a threshold of 5, the LLM achieves an F1 score of 0.98 and 0.59 on the XLCoST and the CodeNet datasets respectively.
@@
[RQ-1.3] ☛ Our experiments on the output requirements (Yes/No vs Similarity scores) underline that both approaches lead to similar performance scores. This suggest that the LLMs indeed considers the clone detection problem as a similarity computation task.
@@
XLCoST CodeNet LLM Clone type Recall Precision F1-score Recall Precision F1-score Falcon-Instruct-7B Non-clone 0.15 0.44 0.22 0.07 0.50 0.12 Clone 0.81 0.49 0.61 0.93 0.50 0.65 StarChat-Beta-16B Non-clone 0.87 0.71 0.78 0.91 0.55 0.69 Clone 0.65 0.84 0.73 0.23 0.70 0.35 LLaMA2-Chat-7B Non-clone 0.74 0.87 0.80 0.65 0.65 0.65 Clone 0.89 0.78 0.83 0.64 0.65 0.64 GPT-3.5-Turbo Non-clone 0.94 0.98 0.96 0.91 0.70 0.79 Clone 0.98 0.94 0.96 0.60 0.87 0.71
4.2. [RQ2] LLMs’ Reasoning for Cross-Lingual Code Clone Detection
Goal. Our investigation of RQ.1 on the performance of LLMS for the cross-lingual code clone detection task has revealed, based on a manual analysis of some sample cases, that the code pairs are often obviously (for a human) a clone, and yet the LLM fails to identify it. In other cases, the pair is obviously (for a human) not a clone and yet the LLM concludes that it is. We thus now aim to further study the extent to which the LLMs understand the task. Based on our analysis, notably of whether the LLMs understand what ”clone” means in a cross-lingual setting, we propose an improved version of the Zero-shot prompt to validate that LLMs performance can be improved with a hint on the definition of clone.
Results. Our qualitative review of LLMs’ outputs reveals that Falcon-Instruct-7B and LLAMA2-Chat-7B consider, most of the time, any input pair of code fragments as clones, leading to an excessive amount of false positives. Conversely, Starchat-$\beta$ considers most of the pairs as non-clones, yielding a significant number of false negatives.
The behavior of Starchat-$\beta$ is explained by the fact the reasoning of the LLM completely breaks down for the task of cross-lingual code clone detection. The LLM even goes far as to over-confidently state that two code snippets of different programming languages cannot be clones. Similar breakdowns in LLM reasoning have recently been documented in the literature (Nezhurina et al., 2024). Our experiments reveal for future research development in LLMs that cross-lingual code clone detection is a relevant and easy-to-assess task for checking potential reasoning breakdown of LLMs. Unfortunately, Starchat-$\beta$ is not the only LLM with a breakdown in reasoning: Similar situations are also apparent in several outputs of GPT-3.5-turbo.
Table 4 presents the performance achieved by the various LLMs based on the improved prompt. The performance improvement (compared to results when using a prompt that simply asks about ”clones” - cf. Table 3) is substantial. In particular, for Starchat-$\beta$ and LLAMA2-Chat-7B, the new prompt helps the LLM to exhibit an improvement between 27 and 48 percentage points in terms of F1 score.
[RQ-2] ☛ LLMs consistently are not fully aware of the notion of code clones in a cross-lingual setting. When prompted to adhere a definition of cloning as a measurement of similarity taking into account overall structure and logic, LLMs are able to demonstrate substantial success in the task of cross-lingual code clone detection. On the XLCoST dataset, where cross-lingual clones are known (cf. RQ1) to be easier to detect, the improved prompt enables the LLaMA2-Chat-7B model to raise its F1 performance score from 0 to 0.80 for non-clones and 0.67 to 0.83 for clones, reducing the gap from GPT-3.5-Turbo by 14.5 percentage points.
@@
Goal. In the studied task, a clone pair involves two fragments from two distinct programming languages. Previous research questions focused on highlight the overall performance of LLMs on randomly built pairs. We now attempt to check whether the specific combinations of programming languages may affect the performance of the LLM in identifying clones. In particular, we investigate whether the syntactic similarity of programming languages has an influence on the effectiveness of the LLMs.
Experiments. Some programming languages share common core concepts, such as for controlling program flow (e.g., loops, conditionals, etc.) or storing data (e.g., variables). In terms of Syntax, we note for example that Java’s syntax is most like C++ and C. In fact, Java borrows heavily from C++ syntax, and therefore their programs may present similar structures. Similarly, languages like C# (pronounced C-Sharp) share similarities with Java because they are object-oriented languages with similar features like classes, objects, inheritance, and encapsulation. To better highlight the potential combinations, we focus on the case of Java, and consider sets of pairs where the second programming language is varied across C, C#, C++, Go, Ocalm, PHP, Python, Ruby, and Rust. We then disaggregate the performance metrics by programming language pair, focusing on experimental results with the CodeNet dataset from RQ-1 (cf. Section 4.1). Given that we investigate the performance across the various prompt designs, we focus on the GPT-3.5-Turbo LLM, for which all prompt designs successfully returned answers.
Results. Table 5 shows all the results across various prompts and by programming language pair. We note that the performance difference is at least of $\sim$10 percentage points in terms of F1 score when the LLM is applied to fragments of Java-C# language pair vs fragments of Java-Python language pair. C# is indeed very close to Java in terms of syntax and core concepts, while Python employs different programming paradigms.
@@
C 0.65 0.76 0.64 0.65 0.59 0.61 0.48
C# 0.62 0.80 0.68 0.69 0.65 0.66 0.54
C++ 0.64 0.79 0.68 0.67 0.60 0.60 0.46
Go 0.57 0.71 0.59 0.60 0.57 0.53 0.47
Ocalm 0.56 0.74 0.65 0.69 0.55 0.60 0.46
PHP 0.62 0.78 0.69 0.73 0.66 0.66 0.56
Python 0.52 0.70 0.56 0.58 0.51 0.51 0.45
Ruby 0.57 0.74 0.62 0.64 0.57 0.56 0.49
Rust 0.64 0.75 0.65 0.66 0.55 0.59 0.46
[RQ-3] ☛ For the task of cross-lingual code clone detection, LLMs, when instructed with simple zero-shot prompts that do not clarify the definition of clones, see their performance influenced by the similarity of the languages of the code fragments in a pair. However, when the prompt explores the reasoning potential of the LLM, the differences in the programming language becomes less critical.
(a) The performance of Text-Embedding-Ada-002 mixed with cosine similarity on XLCoST dataset
(b) The performance of Text-Embedding-Ada-002 mixed with cosine similarity on CodeNet dataset
Figure 3. In (a) and (b), we first measure the performance of Text-Embedding-Ada-002 mixed with cosine similarity at different thresholds for each dataset. We measure the best F1-score, and for the XLCoST dataset the threshold is set at 0.87 and for codeNet at 0.79
@@
4.4. [RQ4] Traditional classification vs. LLMs
Goal. Code clone detection has been studied in the literature as a binary classification problem. Various techniques leveraging machine learning has then been adopted and showed promising results. We compare LLMs against implementations of such techniques. Since we aim to assess the added value of LLMs (beyond the representations), we build on the embedding model that is provided alongside the LLM by OpenAI.
Experiments. We rely on the identified embedding model (Text-Embedding-Ada-002 - cf. Section 3.2) to generate the representations of code fragments in our constructed benchmark. We conduct two separate experiments:
@@
    straightforward similarity computation: we use cosine similarity  (Pedregosa et al., 2011) to compute the distance between the embeddings of two code fragments in a given pair. We can identify an optimal threshold for code clone identification by considering varying threshold values.

### v2

@@ 3.5. Methodology
We overview the experimental methodology that is employed for assessing LLMs as well as baselines, and the classification models for the task of cross-lingual code clone detection. Figure 1 illustrates the process, highlighting the input datasets, the models as well as the expected outputs.
@@
Step 2 : Analyze the following functions of two code snippets and determine if they are code clones, regardless of the programming language. The function of the first code snippet is {step 1 output} and the function of the second is {step 1 output}. Please answer ‘yes’ if the code snippets are clones, regardless of the programming language, or ‘no’ if they are not. yes/no
@@
Cross-lingual code clone detection as a Classification task.
  • Text-embedding-3-large Model.
We investigate, under the same conditions as LLMs, the potential for traditional classification mechanisms based on machine learning models. To that end, we follow the steps used by Keller et al. (Keller et al., 2022) in their learning-based approach for code clone detection. In their approach a classifier is trained based on sample sets of clone and non-clone pairs where each pair is represented by the absolute difference between the embedding vectors of each code fragment in the pair. Compared to other concatenation techniques that were assessed, their work conclude that differencing achieved the best results. In our case, however, the representations are yielded Embedding models that are provided alongside the LLM infrastructure of OpenAI. We use Text-embedding-3-large  [^6] https://openai.com/index/new-and-improved-embedding-model/ embedding model, which was trained on massive code and text data to yield representations that maximally capture semantics and can serve for similarity analysis. Finally, we train a binary classifier based on the yielded representations and using two basic learners: k Nearest Neighbors (k-NN) and Support Vector Machines (SVM).
Beyond the learning-based classification approach, we also investigate a straightforward approach where we compute the similarity between two fragments in a cross-lingual pair represented in a single representation space. To that end, we apply the cosine similarity metric on the Ada embeddings and check for the optimal threshold value to achieve the highest performance in terms of clone pair identification.
  • Baselines Models.
We fine-tune the selected baseline models to optimize their performance for the task. These fine-tuned baselines are then systematically compared to the large language models (LLMs) to evaluate their relative effectiveness. This comparison allows us to assess how traditional models perform compared to the LLMs across the same task, providing insights into the strengths and limitations of each approach.
@@ 4. Evaluation
Simple prompt Non-clone 0.99 0.97 0.98 1.00 0.56 0.72
Clone 0.97 0.99 0.98 0.21 1.00 0.34
Khajezade et al.  (Khajezade et al., 2024) Non-clone - - - 1.00 0.56 0.72
Clone - - - 0.21 1.00 0.35
Similar line Non-clone 1.00 0.98 0.99 1.00 0.60 0.75
Clone 0.98 1.00 0.99 0.33 0.99 0.49
Reasoning Non-clone 0.98 0.99 0.98 0.99 0.71 0.82
Clone 0.99 0.98 0.98 0.59 0.98 0.74
Integrate Non-clone 0.99 0.90 0.94 0.99 0.60 0.74
Clone 0.89 0.99 0.93 0.32 0.98 0.49
Separate code Non-clone 1.00 0.98 0.99 1.00 0.58 0.73
Clone 0.98 1.00 0.99 0.27 1.00 0.42
Separate explanations Non-clone 0.98 0.99 0.98 0.99 0.73 0.84
Clone 0.99 0.98 0.98 0.64 0.99 0.78
@@
XLCoST CodeNet LLM Clone type Recall Precision F1-score Recall Precision F1-score Falcon-Instruct-7B Non-clone 0.15 0.51 0.23 0.11 0.48 0.17 Clone 0.86 0.50 0.63 0.89 0.50 0.64 StarChat-Beta-16B Non-clone 0.62 0.47 0.53 0.76 0.49 0.60 Clone 0.30 0.45 0.36 0.22 0.48 0.30 LLaMA2-Chat-7B Non-clone 0.01 1.00 0.01 0.04 0.68 0.08 Clone 1.00 0.50 0.67 0.98 0.51 0.67 Starcoder2-15b-instruct Non-clone 1.00 0.5 0.67 1.00 0.5 0.67 Clone 0.01 1.00 0.03 0.02 0.96 0.03 GPT-3.5-Turbo Non-clone 0.99 0.97 0.98 1.00 0.56 0.72 Clone 0.97 0.99 0.98 0.21 1.00 0.34
@@
Goal. We aim to fill a gap in the literature with a comprehensive analysis of the effectiveness of LLMs on the task of cross-lingual code clone detection. In particular, we investigate the impact of state-of-the-art prompt engineering strategy on the performance of the LLMs: we thus experiment with Zero-shot and Chain-of-Thought (CoT)-based prompts. Depending on the desired output formats, the designed prompts fall into two categories:
@@
Experiments. We evaluated the five LLMs considered in this study (cf. Section 3.2), namely GPT-3.5-Turbo, Starchat-$\beta$-16B, Falcon-7B-Instruct, LLAMA2-Chat-7B, and, Starcoder2-15b-instruct on the constructed subsets Section 3.3) from XLCoST and CodeNet. We use the ChatGPT API to prompt the GPT-3.5-Turbo LLM and parse the outputs to compute the precision, recall, and F1 scores finally. For prompts leading to outputs with a similarity score, we analyzed the LLMs’ outputs across a range of threshold values for code clone decisions. Our experiments included a comparison of our prompts with a prior proposal for cross-language code clone detection, evaluated on our dataset.
@@
Results. Our experiments confirmed that many LLMs actually fail to support the Chain-of-Thought prompting strategy, confirming a claim by Dou et al. (Dou et al., 2023). Consequently, to analyze the performance of LLMs across different prompts, we focus on GPT-3.5-Turbo, which produced results when instructed with all our designed prompts. Table 2 summarizes the results for this LLM across all the ”yes/no” output prompts. On the XLCoST dataset, the CoT-based ”Separate explanations” prompt enables the model to achieve the highest F1 score (98% for non-clones and clones). On the CodeNet dataset, the same prompt achieves 84% and 78% respectively for non-clones and clones. We can observe that the performance on the CodeNet dataset is lower than that of the XLCoST dataset. This discrepancy in performance is due to the difference in complexity between the two datasets. Indeed, XLCoST is mainly composed of source code implementing simple problems such as “Compute modulus division by a power” while CodeNet is built based on coding challenges of higher complexity such as “Given a string, find the minimum number of swaps needed to rearrange the characters into a palindrome”. Furthermore, some LLMs exhibit a clear difference in their ability to detect code clones compared to non-clones. For instance, on the CodeNet dataset, the GPT-3.5-turbo model shows a 35 percentage point variation in the f1 score (72% on non-clone vs. 37% on clones). On the same dataset, we also observe a high variation of 47 for the ‘Falcon-Instruct-7B’ (17% on non-clones vs. 64% on clones) model. Manual analysis revealed a tendency for this model to classify nearly all input code snippet pairs as clones, resulting in an inflated number of false positives.
@@
[RQ-1.1] ☛ In general, for simple and recurrent programming tasks, LLMs show a high ability to detect cross-lingual clones (e.g., XLCoST samples). However, when applied to code samples of challenging programming tasks (e.g., CodeNet), LLMs performance drop by ${}20\%$ on average in terms of F1 score. This result is insightful as it calls for a different perspective in benchmarking cross-lingual code clone detection approaches. Indeed, while XLCost has been built for such purpose, CodeNet appears to offer a more realistic setting, in terms of complexity for assessment.
@@
The ”Simple prompt” exhibits a relatively consistent performance across both datasets, suggesting its potential for broader applicability. A detailed performance comparison of all five LLMs using this prompt is presented in Table 3. Notably, the GPT-3.5-turbo model outperformed other LLMs in this analysis. Also, our experiments demonstrate that our proposed simple prompt achieved comparable results to the previously suggested one.
@@
[RQ-1.2] ☛ Compared to Zero-Short prompting, which is mainly used in the literature, Chain of Thought prompting improves the performance of LLMs by 1-28 percentage points in terms of overall F1-score. It is further noteworthy that GPT-3.5-Turbo substantially outperforms the other studied LLMs in the task of cross-lingual code clone detection.
@@
In Figure 2, we present the result of the GPT-3.5-turbo LLM with the ”Code Similarity” prompt, designed to generate a similarity score between code snippets (0-10 scale) The results show that with a threshold of 5, the LLM achieves an overall F1 score of 0.98 and 0.78 on the XLCoST and the CodeNet datasets, respectively.
@@
[RQ-1.3] ☛ Our experiments on the output requirements (Yes/No vs Similarity scores) underline that both approaches lead to similar performance scores. This suggests that the LLMs indeed consider the clone detection problem as a similarity computation task.
4.2. [RQ2] LLMs’ Reasoning for Cross-Lingual Code Clone Detection
Goal. Our investigation of RQ.1 on the performance of LLMS for the cross-lingual code clone detection task has revealed, based on a manual analysis of some sample cases, that the code pairs are often obviously (for a human) a clone, and yet the LLM fails to identify it. In other cases, the pair is obviously (for a human) not a clone, and yet the LLM concludes that it is. We thus now aim to further study the extent to which the LLMs understand the task. Based on our analysis, notably of whether the LLMs understand what ”clone” means in a cross-lingual setting, we propose an improved version of the Zero-shot prompt to validate that LLMs performance can be improved with a hint on the definition of clone.
Results. Our qualitative review of LLMs’ outputs reveals that Falcon-Instruct-7B and LLAMA2-Chat-7B consider, most of the time, any input pair of code fragments as clones, leading to an excessive amount of false positives. Conversely, Starcoder2-15b-instruct and Starchat-$\beta$ considers most of the pairs as non-clones, yielding a significant number of false negatives.
The behavior of Starcoder2-15b-instruct and Starchat-$\beta$ are explained by the fact the reasoning of the LLM completely breaks down for the task of cross-lingual code clone detection. The LLM even goes far as to over-confidently state that two code snippets of different programming languages cannot be clones. Similar breakdowns in LLM reasoning have recently been documented in the literature (Nezhurina et al., 2024). Our experiments reveal for future research development in LLMs that cross-lingual code clone detection is a relevant and easy-to-assess task for checking potential reasoning breakdown of LLMs. Unfortunately, Starcoder2-15b-instruct and Starchat-$\beta$ are not the only LLM with a breakdown in reasoning: Similar situations are also apparent in several outputs of GPT-3.5-turbo.
Table 4 presents the performance achieved by the various LLMs based on the improved prompt. The performance improvement (compared to results when using a prompt that simply asks about ”clones” - cf. Table 3) is substantial. In particular, for Starchat-$\beta$ and LLAMA2-Chat-7B, the new prompt helps the LLM to exhibit an improvement between 7 and 48 percentage points in terms of F1 score.
[RQ-2] ☛ LLMs consistently are not fully aware of the notion of code clones in a cross-lingual setting. When prompted to adhere a definition of cloning as a measurement of similarity taking into account overall structure and logic, LLMs are able to demonstrate substantial success in the task of cross-lingual code clone detection. On the XLCoST dataset, where cross-lingual clones are known (cf. RQ1) to be easier to detect, the improved prompt enables the LLaMA2-Chat-7B model to raise its F1 performance score from 0.01 to 0.80 for non-clones and 0.67 to 0.83 for clones, reducing the gap from GPT-3.5-Turbo by 14.5 percentage points.
@@
XLCoST CodeNet LLM Clone type Recall Precision F1-score Recall Precision F1-score Falcon-Instruct-7B Non-clone 0.23 0.54 0.32 0.15 0.51 0.23 Clone 0.80 0.51 0.64 0.86 0.50 0.63 StarChat-Beta-16B Non-clone 0.66 0.60 0.63 0.75 0.52 0.61 Clone 0.57 0.64 0.60 0.34 0.59 0.43 Starcoder2-15b-instruct Non-clone 0.96 0.58 0.73 0.92 0.52 0.67 Clone 0.31 0.89 0.46 0.17 0.67 0.27 LLaMA2-Chat-7B Non-clone 0.75 0.87 0.81 0.78 0.57 0.66 Clone 0.89 0.78 0.83 0.40 0.65 0.50 GPT-3.5-Turbo Non-clone 0.98 0.99 0.98 1.00 0.75 0.85 Clone 0.99 0.98 0.98 0.66 1.00 0.79
@@
Goal. In the studied task, a clone pair involves two fragments from two distinct programming languages. Previous research questions focused on highlight the overall performance of LLMs on built pairs. We now attempt to check whether the specific combinations of programming languages may affect the performance of the LLM in identifying clones. In particular, we investigate whether the syntactic similarity of programming languages has an influence on the effectiveness of the LLMs.
Experiments. Some programming languages share common core concepts, such as for controlling program flow (e.g., loops, conditionals, etc.) or storing data (e.g., variables). In terms of Syntax, we note for example that Java’s syntax is most like C++ and C. In fact, Java borrows heavily from C++ syntax, and therefore their programs may present similar structures. Similarly, languages like C# (pronounced C-Sharp) share similarities with Java because they are object-oriented languages with similar features like classes, objects, inheritance, and encapsulation. To better highlight the potential combinations, we focus on the case of Java, and consider sets of pairs where the second programming language is varied across C, C#, C++, Go, Ocalm, PHP, Python, Ruby, Rust, and JavaScript. We then disaggregate the performance metrics by programming language pair, focusing on experimental results with the CodeNet dataset from RQ-1 (cf. Section 4.1). Given that we investigate the performance across the various prompt designs, we focus on the GPT-3.5-Turbo LLM, for which all prompt designs successfully returned answers.
Results. Table 5 shows all the results across various prompts and by programming language pair. We note that the performance difference is at least of $\sim$10 percentage points in terms of F1 score when the LLM is applied to fragments of Java-C language pair vs fragments of Java-Ocaml language pair. C is indeed very close to Java in terms of syntax and core concepts, while Ocaml employs different programming paradigms.
@@
C 0.68 0.85 0.70 0.83 0.71 0.84 0.60
C# 0.54 0.86 0.65 0.81 0.66 0.82 0.60
C++ 0.59 0.81 0.66 0.81 0.67 0.85 0.56
Go 0.46 0.81 0.57 0.74 0.61 0.82 0.54
Ocaml 0.48 0.81 0.59 0.73 0.54 0.75 0.56
PHP 0.50 0.78 0.62 0.80 0.64 0.82 0.59
Python 0.52 0.80 0.59 0.76 0.61 0.80 0.59
Ruby 0.49 0.81 0.59 0.78 0.52 0.76 0.57
Rust 0.47 0.86 0.57 0.75 0.57 0.80 0.55
JavaScript 0.53 0.85 0.63 0.80 0.63 0.82 0.56
[RQ-3] ☛ For the task of cross-lingual code clone detection, LLMs, when instructed with simple zero-shot prompts that do not clarify the definition of clones, see their performance influenced by the similarity of the languages of the code fragments in a pair. However, when the prompt explores the reasoning potential of the LLM, the differences in the programming language become less critical.
(a) The performance of Text-embedding-3-large mixed with cosine similarity on XLCoST dataset
(b) The performance of Text-embedding-3-large mixed with cosine similarity on CodeNet dataset
Figure 3. In (a) and (b), we first measure the performance of Text-embedding-3-large mixed with cosine similarity at different thresholds for each dataset. Using a threshold of 0.5, we achieved an F1-score of 0.98 on the XLCOST dataset and 0.74 on the Codenet dataset.
@@
4.4. [RQ4] Binary classifier vs. Baselines vs. LLMS
Goal. The primary objective of this research is to evaluate the performance of Embedding models in detecting cross-lingual code clones and to compare their effectiveness against large language models (LLMs). Code clone detection has been studied in the literature as a binary classification problem. Various techniques leveraging machine learning has then been adopted and showed promising results. We compare LLMs against implementations of such techniques. Since we aim to assess the added value of LLMs (beyond the representations), we build on the embedding model that is provided alongside the LLM by OpenAI.
Experiments. We rely on the identified embedding model (Text-embedding-3-large - cf. Section 3.2) to generate the representations of code fragments in our constructed benchmark. We conduct three separate experiments:
@@
    straightforward similarity computation: we use cosine similarity  (Pedregosa et al., 2011) to compute the distance between the embeddings of two code fragments in a given pair. We can identify an optimal threshold for code clone identification by considering varying threshold values. In this study, the threshold value is set to 5 for all evaluations.